OpenStudy (anonymous):

Determine the amplitude, period, and phase shift. y=1/2cos(3x+pi/2)

OpenStudy (jdoe0001):

well... I should say \(\qquad \qquad \qquad \qquad \textit{function transformations} \\ \quad \\ % function transformations for trigonometric functions \begin{array}{rllll} % left side templates f(x)=&{\color{purple}{ A}}cos({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ \end{array}\qquad \begin{array}{llll} % right side info \bullet \textit{ stretches or shrinks horizontally by amplitude } |{\color{purple}{ A}}|\\\ \bullet \textit{ horizontal shift by }\frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is negative, to the right}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is positive, to the left}\\ \bullet \textit{vertical shift by }{\color{green}{ D}}\\ \qquad if\ {\color{green}{ D}}\textit{ is negative, downwards}\\ \qquad if\ {\color{green}{ D}}\textit{ is positive, upwards}\\ \bullet \textit{function period}\\ \qquad \frac{2\pi }{{\color{blue}{ B}}}\ for\ cos(\theta),\ sin(\theta),\ sec(\theta),\ csc(\theta)\\ \qquad \frac{\pi }{{\color{blue}{ B}}}\ for\ tan(\theta),\ cot(\theta) \end{array}\) what do you get from that then?

OpenStudy (jdoe0001):

hmmm anyway... \(\qquad \qquad \qquad \qquad \textit{function transformations} \\ \quad \\ % function transformations for trigonometric functions \begin{array}{rllll} % left side templates f(x)=&{\color{purple}{ \frac{1}{2}}}cos({\color{blue}{ 3}}x+{\color{red}{ \frac{\pi }{2}}})+{\color{green}{ 0}} \\ \quad \\ \end{array}\qquad \begin{array}{llll} % right side info \bullet \textit{ stretches or shrinks}\\\quad \textit{ horizontally by amplitude } |{\color{purple}{ A}}|\\\ \bullet \textit{ horizontal shift by }\frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is negative, to the right}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is positive, to the left}\\ \bullet \textit{vertical shift by }{\color{green}{ D}}\\ \qquad if\ {\color{green}{ D}}\textit{ is negative, downwards}\\ \qquad if\ {\color{green}{ D}}\textit{ is positive, upwards}\\ \bullet \textit{function period}\\ \qquad \frac{2\pi }{{\color{blue}{ B}}}\ for\ cos(\theta),\ sin(\theta),\ sec(\theta),\ csc(\theta)\\ \qquad \frac{\pi }{{\color{blue}{ B}}}\ for\ tan(\theta),\ cot(\theta) \end{array}\) see what you get